OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@splatter6

OpenStudy (anonymous):

y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@ichibanstunna

OpenStudy (anonymous):

C

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np